# EMT: A Visual Multi-Task Benchmark Dataset for Autonomous Driving in the Arab Gulf Region

### v1

@@ IV Evaluation
TABLE IV: Off-shelf vs Fine tuned YOLOX-L detection results (the best results are highlighted in bold)
Detector Object Class Evaluation Metrics
Precision↑ Recall↑ F1Score ↑ FP↓ FN↓
YoloX Off-shelf Vehicle 90.14 23.59 37.39 2,669 79,002
Pedestrian 65.59 10.33 17.84 255 4,221
Motorbike 96.15 4.61 8.79 4 2,071
Cyclist 50.00 0.65 1.28 1 153
overall 89.50 22.62 36.11 2,929 85,447
YoloX fine-tunned Vehicle 85.53 66.87 75.06 11,692 34,257
Pedestrian 82.76 73.13 77.64 717 1,265
Motorbike 93.12 56.70 70.48 91 940
Cyclist 93.63 95.45 94.53 10 7
overall 85.53 66.97 75.12 12,510 36,469
TABLE V: Tracking Evaluation Results for six Detector-Tracker Settings (the best results for fine-tuned setting between two trackers are highlighted in bold)
Det. Tracker Object Class Evaluation Metrics
MOTA↑ HOTA↑ IDF1↑ FP↓ FN↓ IDs↓
Ground Truth BoT-SORT Vehicle 92.0 73.9 74.8 117,353 147,168 12,992
Pedestrian 84.9 70.9 76.8 4,158 6,674 1,156
Motorbike 86.8 77.1 79.6 1,542 2,800 200
Cyclist 93.9 77.02 89.0 50 79 7
ByteTrack Vehicle 85.4 66.9 72.5 124,039 162,065 14,044
Pedestrian 70.6 59.8 68.0 4,676 9,493 1,742
Motorbike 83.0 71.4 76.8 1,723 3,180 208
Cyclist 73.9 60.3 75.8 61 194 16
YoloX Off-shelf BoT-SORT Vehicle 20.2 25.9 29.4 20,924 443,214 1,077
Pedestrian 6.6 14.1 14.0 757 22,656 53
Motorbike 3.7 8.3 5.9 135 10,944 10
Cyclist 0 6.5 6.4 40 573 1
ByteTrack Vehicle 19.8 25.4 29.6 25,507 441,511 1,757
Pedestrian 6.9 13.7 13.8 721 22,689 67
Motorbike 3.7 8.5 5.9 157 10,940 13
Cyclist 0 3.6 3.5 27 583 0
YoloX fine-tunned BoT-SORT Vehicle 72.7 57.6 65.8 130,451 211,305 11,009
Pedestrian 58.7 50.4 63.0 4,909 11,011 960
Motorbike 63.0 58.1 68.2 1,255 4,804 127
Cyclist 64.8 54.5 69.9 74 235 6
ByteTrack Vehicle 67.4 55.0 64.7 132,892 217,837 12,055
Pedestrian 48.5 44.8 55.4 5,175 13,171 1,361
Motorbike 60.7 56.2 65.9 1,377 5,063 113
Cyclist 45.8 40.0 56.3 57 339 14
@@ IV-A1 Trackers
  • BoT-SORT [20] is a multi-object tracker that enhances motion-based tracking through camera-motion compensation and an improved Kalman filter state representation. Like ByteTrack, it employs a two-level association strategy. The tracker employs a cascade matching approach that prioritizes motion information for data association, with an optional ReID extension that adds appearance features as a secondary matching cue.
@@ IV-A2 Evaluation Protocol
We evaluate the trackers using three different detector configurations:
  • ground-truth detections assuming the presence of perfect detector, providing an upper bound on performance;
  • off-the-shelf YOLO detector, representing readily available solutions;
  • fine-tuned YOLO detector, optimized for our specific use case.
Each detector configuration serves a distinct analytical purpose. Ground-truth detections allow us to evaluate the tracker’s association algorithm under ideal conditions, isolating its core tracking capabilities from detection errors. The off-the-shelf YOLO detector helps assess tracker robustness under challenging conditions with missing and erroneous detections. Finally, the fine-tuned YOLO detector represents a more realistic production scenario, where the detector is optimized for the specific domain but still maintains some inherent errors. We conduct the evaluation using F1-score, Identity Switches and Higher Order Tracking Accuracy (HOTA):
@@
For evaluation purposes, we group detection classes into four superclasses: pedestrian, motorbike, cyclist, and vehicle. The vehicle superclass encompasses all other classes in the dataset, including cars, buses, small motorized vehicles, medium and large vehicles, and emergency vehicles. Tracking performance metrics are broke down for each superclass to provide a detailed analysis of tracker behavior across different object categories. The evaluation was done on the full dataset, i.e., including train and test splits.
@@ IV-A4 Results
After fine-tuning YOLOX-L on our dataset, we evaluated both the fine-tuned and off-the-shelf models as detection backends for tracking. Table IV presents the per-class detection results. All evaluations used a detection confidence threshold of 0.4 and an NMS threshold of 0.5. While NMS is typically set higher, the densely packed objects and frequent bounding box overlaps in our dataset required a lower threshold to prevent excessive suppression. The fine-tuned YOLOX-L model significantly improved recall across all object classes, leading to higher F1 scores, particularly for pedestrians, motorbikes, and cyclists. The F1 score increased by approximately 2.5x across classes compared to the off-the-shelf model. While for vehicles, the precision dropped from 90.14 to 85.53, this trade-off resulted in a substantial reduction in false negatives (FN), cutting overall missed detections from 85,447 to 36,469. This improvement is crucial for tracking, where recall is essential for maintaining track consistency. Error analysis highlights a key limitation of the off-the-shelf model. Its high precision comes at the cost of poor recall, leading to frequent missed detections. This performance gap is particularly pronounced in challenging scenarios, such as high object density and adverse conditions like rainy nights. Fine-tuning on the train split enabled better detection robustness, especially for underrepresented classes.
Table V presents tracking results across three detection settings and two tracking models, yielding six result combinations. BoT-SORT consistently outperforms ByteTrack in HOTA and MOTA across all settings, except for motorbike HOTA and pedestrian MOTA in the off-the-shelf setting. The largest performance gap is observed in the ground-truth setting for cyclists, where BoT-SORT achieves a MOTA improvement of 20 and HOTA improvement of 16.7 over ByteTrack. The low false negatives (FN) and high false positives (FP) in the off-the-shelf setting stem from the detector’s low recall, making it overly permissive and generating excessive detections with poor precision. For instance, in ByteTrack, FP for vehicles drops from 124,039 (ground-truth) to 20,924 (off-the-shelf), but FN increases more than 2.5×, from 162,065 to 441,511, confirming that the model struggles with false detections despite detecting more objects.
@@
TABLE VI: Trajectory Predictors Evaluation (Sequential)
@@
TABLE VII: Trajectory Predictors Evaluation (Multimodal, Sequential)
@@
TABLE VIII: Trajectory Predictors Evaluation (Frame-based)
@@ IV-B3 Implementation Details
To construct the mask, given a batch with a variable number of agents $n$ (up to max_num_nodes), the mask is computed as:
$M_{i}=\begin{cases}1,&\text{if }i\leq n,\\ 0,&\text{otherwise}.\end{cases}$
This mask is applied during loss computation to prevent gradient updates on padded entries. Edges in the adjacency matrix are established if the pairwise distance between nodes is below a predefined threshold, which is set to 100. To ensure stable training and effective message passing, the adjacency matrix is symmetrically normalized:
@@ IV-B4 Results
Table VI presents the evaluation results of LSTM, Transformer, and Transformer coupled with GMM on prediction datasets with varying prediction horizons, observed trajectory lengths, and sliding window settings. The best results for each column are highlighted in bold, indicating which model and sliding window configuration performed best. For longer prediction horizons (i.e., 20, 30, and 60), the Transformer coupled with GMM outperforms the other models in terms of final displacement error (FDE). In contrast, for an observed trajectory of 10 and a prediction horizon of 1, LSTM achieves the best performance. The first two settings in the table (10/10 and 20/10) test the hypothesis that a longer observed trajectory improves prediction accuracy. This hypothesis is supported by the results, as using an observed trajectory of 20 achieves an ADE of 9.20 compared to 10.49 for an observed trajectory of 10. Similarly, the FDE improves from 19.63 to 16.84. A similar trend is observed in the third and fourth columns for a prediction horizon of 20 with observed trajectories of 10 and 20. The best settings demonstrate an FDE improvement of 5.36 and an ADE improvement of 2.04 when increasing the observed trajectory length from 10 to 20.
Table VII presents the results of the multimodal output evaluation, with the best results highlighted in bold across different sliding window settings. The model performed best when trained using a sliding window of 1. Compared to the unimodal setting, the multimodal approach demonstrates a clear advantage. For instance, in the 20/60 split, the best ADE in the unimodal setting is 44.94, whereas the multimodal model achieves 32.30. Similarly, FDE improves from 93.28 (unimodal) to 62.23 (multimodal), representing a reduction of over 30. In the simplest setting (10/10), the multimodal model outperforms LSTM, achieving a gain of 3.3 in ADE and 7.6 in FDE. Further, Table VIII presents the results of frame-based evaluation for prediction horizons of 10 and 20. To better reflect real-world applications, objects without a sufficiently long past trajectory, matching the model’s required observation length, are padded with zeros to simulate newly detected objects entering the scene. For a prediction horizon of 10, LSTM demonstrates the most robust performance across all test videos, except for the third sample, where the Transformer achieves a better score by 0.48 in ADE and 0.35 in FDE. In the case of a longer prediction horizon, GAT outperforms LSTM in the second and third samples, achieving more tangible improvements, with an average gain of 2.5 in ADE and 7.06 in FDE. Figure 4 illustrates examples of predicted trajectories, where LSTM-generated predictions in the 10/10 setting are shown with green arrows, while ground truth trajectories are depicted in red. The predictions accurately capture the agent’s heading but show lower accuracy in precise location estimation, particularly struggling to adapt to varying agent speeds.
@@ IV-C Intention Prediction
TABLE IX: Intention Prediction Evaluation (I - vanilla setting and II - autoregressive mode).
@@ IV-C4 Results
Table IX presents the results of the conducted experiments, detailing the performance across both model settings, intention classes, and past trajectory settings. The highest performance is achieved by the vanilla model configuration, surpassing the autoregressive setting by nearly 2% in F1-score across all timestamps. The results indicate that when only relative displacement is utilized for prediction, the model fails to capture temporal dependencies between intentions. Intentions such as ”waiting_to_cross” and ”crossing” should exhibit sequential dependencies. However, as shown in Figure 5, the model incorrectly predicts the ”keep_lane” intention for crossing scenarios. This misclassification occurs due to the similarity between lane-keeping motion patterns and the vehicle-relative displacement of crossing pedestrians. Additionally, since the model does not incorporate object type information or associated intention priors, it is unable to distinguish between different entities.
@@ V Conclusion and Future Work
This paper introduced the Emirates Multi-Task (EMT) dataset, which includes tracking, trajectory prediction, and intention prediction datasets. Each dataset is accompanied by evaluation code and results from conducted experiments. For tracking, the fine-tuned detector significantly improves performance, reinforcing the importance of using region-specific data for effective adaptation of deep learning algorithms in autonomous vehicle applications. This highlights the necessity of fine-tuning models on local data to achieve optimal performance. For trajectory prediction, we evaluate both sequential and interaction-aware models, providing training pipelines for sequential and frame-based learning. The reported benchmark results serve as a reference for future research. For intention prediction, two evaluation settings are provided: cross-validation and train/test splits. We report cross-validation results as a benchmark, while the train/test setting is intended for assessing model generalization and performance on previously unseen intention classes. The EMT dataset is designed to facilitate the evaluation of models on the unique characteristics of the Gulf region, offering a valuable resource for advancing research in autonomous driving.

### v3

@@ IV Evaluation
TABLE V: Off-shelf vs Fine tuned YOLOX-L detection results (the best results are highlighted in bold)
Object Class YoloX Off-shelf YoloX fine-tuned
Precision↑ Recall↑ F1Score ↑ FP↓ FN↓ Precision↑ Recall↑ F1Score ↑ FP↓ FN↓
Vehicle 90.14 23.59 37.39 2,669 79,002 85.53 66.87 75.06 11,692 34,257
Pedestrian 65.59 10.33 17.84 255 4,221 82.76 73.13 77.64 717 1,265
Motorbike 96.15 4.61 8.79 4 2,071 93.12 56.70 70.48 91 940
Cyclist 50.00 0.65 1.28 1 153 93.63 95.45 94.53 10 7
overall 89.50 22.62 36.11 2,929 85,447 85.53 66.97 75.12 12,510 36,469
TABLE VI: Tracking Evaluation Results for six Detector-Tracker Settings (the best results for fine-tuned setting between two trackers are highlighted in bold)
Det. Object Class BoT-SORT ByteTrack
MOTA↑ HOTA↑ IDF1↑ FP↓ FN↓ IDs↓ MOTA↑ HOTA↑ IDF1↑ FP↓ FN↓ IDs↓
Ground Truth Vehicle 92.0 73.9 74.8 117,353 147,168 12,992 85.4 66.9 72.5 124,039 162,065 14,044
Pedestrian 84.9 70.9 76.8 4,158 6,674 1,156 70.6 59.8 68.0 4,676 9,493 1,742
Motorbike 86.8 77.1 79.6 1,542 2,800 200 83.0 71.4 76.8 1,723 3,180 208
Cyclist 93.9 77.02 89.0 50 79 7 73.9 60.3 75.8 61 194 16
YoloX Off-shelf Vehicle 20.2 25.9 29.4 20,924 443,214 1,077 19.8 25.4 29.6 25,507 441,511 1,757
Pedestrian 6.6 14.1 14.0 757 22,656 53 6.9 13.7 13.8 721 22,689 67
Motorbike 3.7 8.3 5.9 135 10,944 10 3.7 8.5 5.9 157 10,940 13
Cyclist 0 6.5 6.4 40 573 1 0 3.6 3.5 27 583 0
YoloX fine-tunned Vehicle 72.7 57.6 65.8 130,451 211,305 11,009 67.4 55.0 64.7 132,892 217,837 12,055
Pedestrian 58.7 50.4 63.0 4,909 11,011 960 48.5 44.8 55.4 5,175 13,171 1,361
Motorbike 63.0 58.1 68.2 1,255 4,804 127 60.7 56.2 65.9 1,377 5,063 113
Cyclist 64.8 54.5 69.9 74 235 6 45.8 40.0 56.3 57 339 14
@@ IV-A1 Trackers
  • BoT-SORT [25] is a multi-object tracker that enhances motion-based tracking through camera-motion compensation and an improved Kalman filter state representation. Like ByteTrack, it employs a two-level association strategy. The tracker employs a cascade matching approach that prioritizes motion information for data association.
@@ IV-A2 Evaluation Protocol
We evaluate the trackers using three different detector configurations: (1) ground-truth detections assuming the presence of perfect detector, providing an upper bound on performance; (2) off-the-shelf YOLO detector, representing readily available solutions; (3) fine-tuned YOLO detector, tuned for our use case. Each detector configuration serves a distinct analytical purpose. Ground-truth detections allow us to evaluate the tracker’s association algorithm under ideal conditions, isolating its core tracking capabilities from detection errors. The off-the-shelf YOLO detector helps assess tracker robustness under challenging conditions with missing and erroneous detections. Finally, the fine-tuned YOLO detector represents a more realistic production scenario, where the detector is optimized for the specific domain but still maintains some inherent errors. We conduct the evaluation using F1-score, Identity Switches and Higher Order Tracking Accuracy (HOTA):
@@
To match YOLO’s classes, we group EMT object classes into four superclasses: pedestrian, motorbike, cyclist, and vehicle. The vehicle superclass encompasses all other classes in the dataset, including cars, buses, small motorized vehicles, medium and large vehicles, and emergency vehicles. Tracking performance metrics are broke down for each superclass to provide a detailed analysis of tracker behavior across different object categories. The evaluation was done on the full dataset, i.e., including train and test splits.
@@ IV-A4 Results
After fine-tuning YOLOX-L on our dataset, we evaluated both the fine-tuned and off-the-shelf models as detection backends for tracking. Table V presents the per-class detection results. All evaluations used a detection confidence threshold of 0.4 and an NMS threshold of 0.5. While NMS is typically set higher, the densely packed objects and frequent bounding box overlaps in our dataset required a lower threshold to prevent excessive suppression. The fine-tuned YOLOX-L model significantly improved recall across all object classes, leading to higher F1 scores, particularly for pedestrians, motorbikes, and cyclists. The F1 score increased by approximately 2.5x across classes compared to the off-the-shelf model. While for vehicles, the precision dropped from 90.14 to 85.53, this trade-off resulted in a substantial reduction in false negatives (FN), reducing overall missed detections from 85,447 to 36,469. This improvement is crucial for tracking, where recall is essential for maintaining track consistency. Error analysis highlights a key limitation of the off-the-shelf model. Its high precision comes at the cost of poor recall, leading to frequent missed detections. This performance gap is particularly pronounced in challenging scenarios, such as high object density and adverse conditions like rainy nights. Fine-tuning on the train split enabled better detection robustness, especially for underrepresented classes.
Table VI presents tracking results across three detection settings and two tracking models, yielding six result combinations. BoT-SORT consistently outperforms ByteTrack in HOTA and MOTA across all settings, except for motorbike HOTA and pedestrian MOTA in the off-the-shelf setting. The largest performance gap is observed in the ground-truth setting for cyclists, where BoT-SORT achieves a MOTA improvement of 20 and HOTA improvement of 16.7 over ByteTrack. The low false negatives (FN) and high false positives (FP) in the off-the-shelf setting stem from the detector’s low recall, making it overly permissive and generating excessive detections with poor precision. For instance, in ByteTrack, FP for vehicles drops from 124,039 (ground-truth) to 20,924 (off-the-shelf), but FN increases more than 2.5×, from 162,065 to 441,511, confirming that the model struggles with false detections despite detecting more objects.
@@
TABLE VII: Trajectory Predictors Evaluation (Sequential)
@@
TABLE VIII: Trajectory Predictors Evaluation (Multimodal, Sequential)
@@
TABLE IX: Trajectory Predictors Evaluation (Frame-based)
@@ IV-B3 Implementation Details
To construct the mask, given a batch with a variable number of agents $n$ (up to max_num_nodes), the mask is computed as $M_{i}=1,\text{if }i\leq n$ and $0,\text{otherwise}$. This mask is applied during loss computation to prevent gradient updates on padded entries. Edges in the adjacency matrix are established if the pairwise distance between nodes is below a predefined threshold, which is set to 100. To ensure stable training and effective message passing, the adjacency matrix is symmetrically normalized:
@@ IV-B4 Results
Table VII presents the evaluation results of LSTM, Transformer, and Transformer coupled with GMM on prediction datasets with varying prediction horizons, observed trajectory lengths, and sliding window settings. The best results for each column are highlighted in bold, indicating which model and sliding window configuration performed best. For longer prediction horizons (i.e., 20, 30, and 60), the Transformer coupled with GMM outperforms the other models in terms of final displacement error (FDE). In contrast, for an observed trajectory of 10 and a prediction horizon of 1, LSTM achieves the best performance. The first two settings in the table (10/10 and 20/10) test the hypothesis that a longer observed trajectory improves prediction accuracy. This hypothesis is supported by the results, as using an observed trajectory of 20 achieves an ADE of 9.20 compared to 10.49 for an observed trajectory of 10. Similarly, the FDE improves from 19.63 to 16.84. A similar trend is observed in the third and fourth columns for a prediction horizon of 20 with observed trajectories of 10 and 20. The best settings demonstrate an FDE improvement of 5.36 and an ADE improvement of 2.04 when increasing the observed trajectory length from 10 to 20.
Table VIII presents the results of the multimodal output evaluation, with the best results highlighted in bold across different sliding window settings. The model performed best when trained using a sliding window of 1. Compared to the unimodal setting, the multimodal approach demonstrates a clear advantage. For instance, in the 20/60 split, the best ADE in the unimodal setting is 44.94, whereas the multimodal model achieves 32.30. Similarly, FDE improves from 93.28 (unimodal) to 62.23 (multimodal), representing a reduction of over 30. In the simplest setting (10/10), the multimodal model outperforms LSTM, achieving a gain of 3.3 in ADE and 7.6 in FDE. Further, Table IX presents the results of frame-based evaluation for prediction horizons of 10 and 20. To better reflect real-world applications, objects without a sufficiently long past trajectory, matching the model’s required observation length, are padded with zeros to simulate newly detected objects entering the scene. For a prediction horizon of 10, LSTM demonstrates the most robust performance across all test videos, except for the third sample, where the Transformer achieves a better score by 0.48 in ADE and 0.35 in FDE. In the case of a longer prediction horizon, GAT outperforms LSTM in the second and third samples, achieving more tangible improvements, with an average gain of 2.5 in ADE and 7.06 in FDE. Figure 4 illustrates examples of predicted trajectories, where LSTM-generated predictions in the 10/10 setting are shown with green arrows, while ground truth trajectories are depicted in red. The predictions accurately capture the agent’s heading but show lower accuracy in precise location estimation, particularly struggling to adapt to varying agent speeds.
@@ IV-C Intention Prediction
TABLE X: Intention Prediction Evaluation (I - vanilla setting and II - autoregressive mode).
@@ IV-C4 Results
Table X presents the results of the conducted experiments, detailing the performance across both model settings, intention classes, and past trajectory settings. The highest performance is achieved by the vanilla model configuration, surpassing the autoregressive setting by nearly 2% in F1-score across all timestamps. The results indicate that when only relative displacement is utilized for prediction, the model fails to capture temporal dependencies between intentions. Intentions such as ”waiting_to_cross” and ”crossing” should exhibit sequential dependencies. However, as shown in Figure 5, the model incorrectly predicts the ”keep_lane” intention for crossing scenarios. This misclassification occurs due to the similarity between lane-keeping motion patterns and the vehicle-relative displacement of crossing pedestrians. Additionally, since the model does not incorporate object type information or associated intention priors, it is unable to distinguish between different entities.
@@ V Conclusion and Future Work
This paper introduced the Emirates Multi-Task (EMT) dataset, designed to support the evaluation of tracking, trajectory prediction, and intention prediction within a unified multi-task framework. On cross-task evaluation, in detection-tracking pipeline, the fine-tuned detector significantly improves performance, reinforcing the importance of using region-specific data for effective adaptation of deep learning algorithms in autonomous vehicle applications. For trajectory prediction, we evaluate both sequential and interaction-aware models, providing training pipelines for sequential and frame-based learning. For intention prediction, two evaluation settings are provided: cross-validation and train/test splits. We report cross-validation results as a benchmark, while the train/test setting is intended for assessing model generalization on previously unseen intention classes. The broader cross-task investigations, as tracking-to-prediction, detection-to-tracking-to-prediction, tracking-to-intention inference, and joint models where trajectory forecasting is conditioned on predicted intentions are left open for future investigation by the research community and datasets users.